# PBEBench: A Multi-Step Programming by Examples Reasoning Benchmark inspired by Historical Linguistics

### v3

@@ 3.1.3 Enforcing Complexity Constraints
We control the complexity of cascades of string rewrite programs $\vec{p}$ using rejection sampling guided by the classifiers developed in Section C. Complexity is balanced along two dimensions: (1) the cascade length $|\vec{p}|$, and (2) the set of relation types present between programs in $\vec{p}$. The second dimension is encoded as a binary category string $c_{p}\in\{0,1\}^{4}$. Each digit corresponds to the binary presence or absence of feeding, bleeding, counter-feeding, and counter-bleeding relations. Instances are balanced across all 16 possible category strings, from $c_{p}=0000$ (no relations present, arbitrary ordering possible) to $c_{p}=1111$ (all relations present, ordering highly constrained).
Computing Instance Complexity. We hypothesize that the difficulty of a PBE problem $\langle\vec{i},\vec{p},\vec{o}\rangle$ is governed by (1) cascade length and (2) the types of relations between program pairs in $\vec{p}$ (relations formulated by phonologists and historical linguists; see Kiparsky (1968; 1971)). One key contribution is a provably correct and automatic classification of these relations (Appendix C). For any ordered pair of programs ($p_{i}$$p_{j}$) in $\vec{p}$, the possible relations are:
  • Feeding (F($p_{i}$, $p_{j}$): $p_{i}$ creates substrings that enable $p_{j}$ to apply.
  • Bleeding (B($p_{i}$, $p_{j}$): $p_{i}$ removes substrings required by $p_{j}$.
@@
Counter-relations need not be stored explicitly, since $CF(p_{i},p_{j})$ and $CB(p_{i},p_{j})$ are implied if $p_{j}$ precedes $p_{i}$ and $F(p_{j},p_{i})$ and $B(p_{j},p_{i})$, respectively. These relations can be visualized using a Directed Acyclic Graph (DAG), as illustrated in Figure 1. We show both the full DAG with all relations and a simplified DAG that indirectly encodes counter-relations via this symmetry.
Rejection Sampling to Control Complexity. To enforce balanced complexity, we use rejection sampling (Section 3.1.2). Each instance is assigned a category string $c_{p}$, and data is generated to approximate balance across the $2^{4}=16$ relation-type categories. We also aim to balance cascade lengths between $L_{min}$ and $L_{max}$, though these constraints can conflict. For example, with $L=2$, randomly sampled program pairs rarely exhibit three or four relation types, making some category strings unattainable. Prior work on sound law induction (Naik et al., 2025a) showed that fine-tuning on randomly sampled programs yielded greater improvements in real-world forward reconstruction than retrieving from attested sound laws, likely due to higher diversity. Following this insight, we retain random sampling (rather than deterministic generation or database retrieval) to promote diversity and avoid contamination. This leads to high rejection rates, often requiring constraint relaxation. We therefore introduce a patience parameter $\tau$, which is the number of sampling steps carried out with both constraints. Once $\tau(=100000)$ is exceeded, we accept any instance satisfying at least one constraint (either $c_{p}$ or cascade length). In early experiments, relation-type balance was prioritized, but later analysis (Appendix G.3, G.4) showed cascade length to be a stronger predictor of difficulty. Accordingly, in subsequent experiments, we relaxed relation-type constraints while maintaining strict control over cascade length. We also study the effect of using different values of $\tau$ on the balance of relation types and efficiency of the data generation process (Appendix G.2).
@@ 3.2 Problem Solver
Prompting Strategy: Once a benchmark snapshot $\mathcal{D}$ is generated, we evaluate each LLM $M$ by prompting it (see appendix E.2) to produce a candidate program cascade $\hat{\vec{p}}=M(\vec{i},\vec{o})$ mapping inputs $\vec{i}$ to outputs $\vec{o}$. The prompt enforces the following constraints: each program must be a Python replace function; both arguments $\alpha_{k}$ and $\beta_{k}$ must be strings with $|\alpha_{k}|,|\beta_{k}|\leq s_{max}$; $\alpha_{k}$ must be non-empty; the cascade may contain at most $L_{max}$ programs (matching the length of the longest ground-truth cascade); and the output must follow a strict Markdown format for reliable extraction. If a predicted program $\hat{p}_{j}$ has more than $L_{max}$ programs, we consider the first $L_{max}$ programs for evaluation. If the predicted program violates any other constraint, it is replaced by an identity program $p^{I}$, which leaves inputs unchanged. Given the predicted cascade $\hat{\vec{p}}$, we compute the predicted outputs as $\hat{\vec{o}}=\hat{\vec{p}}(\vec{i})$. Since we evaluate all kinds of LLMs, including reasoning models that may first produce intermediate outputs and then iteratively refine or improve them by reflection, we account for this by evaluating both the first and last code block produced by each model.
Scaling Solution Search: We explore two approaches to scaling the solution search, inspired by prior work on LLMs for PBE (Li & Ellis, 2024) and general reasoning (Muennighoff et al., 2025): (1) sampling budget and (2) test-time thinking-budget scaling. The first approach gives the LLM multiple attempts to solve a problem by successively sampling solutions $K$ times and selecting any solution consistent with the input-output examples. If no correct solution is found in the $K$ samples, we pick the one with the highest edit similarity reward (defined in Section 4.3). The second approach is inspired by Muennighoff et al. (2025). We constrain the thinking budget to $N$ tokens and scale it by increasing this limit. However, we observed that gpt-oss-120b does not strictly adhere to thinking budget constraint, even when employing strategies such as forcing it to emit chain-of-thought termination tokens (experimental details are available in Appendix F.3), so we ultimately control the overall maximum sequence length.
@@ 4 Experiments
4.1 Benchmark Creation
Using the problem proposer described in Section 3.1, we construct two benchmarks. PBEBench-Lite is a simplified dataset with longer cascades and 5 input–output pairs per instance, synthesized as $\mathcal{D}(n=5,\Sigma=\Sigma_{\text{lite}},L_{min}=2,L_{max}=5,l_{min}=2,l_{max}=6,s_{min}=1,s_{max}=3,D=1008)$, with exactly 63 examples per relation category. The alphabet is restricted $\Sigma_{\text{lite}}=\{a,\dots,k,u,v,w,x,y,z\}$ (lowercase letters excluding lltt). PBEBench is a larger dataset with 50 input-output pairs per instance, synthesized as $\mathcal{D}(n=50,\Sigma,L_{min}=2,L_{max}=20,l_{min}=2,l_{max}=6,s_{min}=1,s_{max}=3,D=1216)$, with 64 instances for each cascade length. Here $\Sigma=\{a,\dots,z,A,\dots,Z\}$ (the full alphabet in both cases). While PBEBench does not reach the complexity of real-world forward reconstruction, where cascades may exceed 200 programs, it remains challenging for the strongest open-source models. For evaluation of advanced closed-source systems such as GPT-5, we additionally synthesize cases with cascades up to length 30.
@@ 4.2 Models Evaluated
We evaluate a broad range of state-of-the-art LLMs spanning multiple model families and design choices. Our evaluation covers models from leading developers including OpenAI (GPT, o-series, gpt-oss), Anthropic (Claude series), Google DeepMind (Gemini), Qwen (Qwen2.5, Qwen3, QwQ), DeepSeek (R1-Distill-Qwen), Mistral (Codestral). The models differ in scale, reasoning specialization (thinking vs. non-thinking), architectural choices (dense vs. MoE), and source availability (open vs. closed). A full list of models with their attributes is provided in Table 2, which demonstrates the breadth of model families and capabilities we consider.
@@ 4.3 Evaluation Metrics
Since a given list of inputs ($\vec{i}$) could be transformed into the outputs ($\vec{o}$) by multiple program cascades $\vec{p}$ we utilize metrics based on functional equivalence. We execute the model-generated solution $\hat{\vec{p}}$ on the inputs, treating them like test cases. We compare the predicted outputs $\hat{\vec{o}}=\hat{\vec{p}}(\vec{i})$ with the ground truth outputs $\vec{o}$ at two levels of granularity: (1) coarse-grained evaluation (Pass@1 or solve rate) and (2) fine-grained evaluation (Edit_Sim or normalized edit similarity). Both metrics do element-wise comparisons on the strings in the output vectors:
Coarse-grained Metric (Pass@1): This metric is Pass@1 from Chen et al. (2021b).
$\mathrm{pass@1}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{i}\in\mathcal{D}}\mathrm{1}_{\hat{\vec{p}}(\vec{i})=\vec{o}}$
Here $\mathrm{1}_{\hat{\vec{p}}(\vec{i})=\vec{o}}$ is an indicator variable which is 1 when $\hat{\vec{p}}(\vec{i})=\vec{o}$} is true and 0 if it is false.
Fine-grained Metric (Edit_Sim): This metric is the same as Reward@1 used by Naik et al. (2025b).
$\mathrm{Edit\_Sim}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{i}\in\mathcal{D}}1-\frac{\mathrm{dist}(\hat{\vec{p}}(\vec{i}),\vec{o})}{\mathrm{dist}(\vec{i},\vec{o})}$
Here $\mathrm{dist}$ denotes the total Levenshtein edit distance summed across the corresponding inputs and outputs. Additionally, we evaluate the proportion of programs an LLM generates that are valid (follow all instructions mentioned in 3.2) which we term as the Valid_Rate.
Model First Code Block Last Code Block
Pass@1 Edit Sim Valid Rate Pass@1 Edit Sim Valid Rate
Codestral-22B 0.0109 -0.0130 0.8280 0.0109 -0.0107 0.8254
Qwen2.5-32B-Instruct 0.0298 0.1232 0.8265 0.0298 0.1252 0.8288
Qwen3-32B 0.0179 0.0964 0.7645 0.0179 0.0964 0.7645
Qwen3-30B-A3B ★◼ 0.2887 0.3360 0.9905 0.2887 0.3360 0.9905
QwQ-32B ★ 0.3591 0.4072 0.9500 0.3601 0.4092 0.9493
Qwen3-32B (with CoT) ★ 0.3938 0.4803 0.9182 0.4187 0.5026 0.9676
gpt-oss-20b ★◼ 0.4058 0.4619 0.9900 0.4058 0.4619 0.9900
gpt-oss-120b ★◼ 0.6250 0.6985 0.9254 0.6250 0.6985 0.9254
Claude-3.5-Sonnet 0.1845 0.4430 0.8208 0.1845 0.4434 0.8209
Claude-3.7-Sonnet 0.2212 0.4825 0.8409 0.2321 0.4996 0.8409
Claude-4 Sonnet 0.2956 0.5832 0.7720 0.2966 0.5870 0.7719
o3-mini ★ 0.5278 0.5954 0.9179 0.5278 0.5954 0.9179
o4-mini ★ 0.6329 0.6907 0.9153 0.6329 0.6907 0.9153
Gemini 2.5 Flash ★ 0.5833 0.6533 0.7911 0.5863 0.6562 0.7901
GPT-5 ★ 0.7242 0.7645 0.9286 0.7242 0.7645 0.9286
@@ 5.1 PBEBench-Lite Performance
Table 1 reports Pass@1, Edit_Sim, and Valid_Rate for all models (Section 4.2) on both the first and last code block from each response (as explained in Section 3.2). Since the last block usually performs as well or better, we use it for instance difficulty analysis. Hyperparameters used for model evaluations are listed in Table 6. Claude-4 Opus (Thinking) was run on only 20% of data due to high API costs (details for costs of experiment runs are added in Table-E.5. Reasoning models, open and closed-source, consistently outperform non-reasoning ones, with a larger gap for open-source models. The top models are gpt-oss-120b (open) and GPT-5 (closed). Factorial analysis of QwQ-32B and GPT-5 shows cascade length as the strongest negative predictor of Pass@1. We further find that feeding and bleeding decrease Pass@1, while counter-feeding and counter-bleeding have no significant effect. (Appendix G.3). Finally, we compare the distributions of cascade lengths and relation types in model predictions to the ground truth using confusion matrices (Appendix G.5, G.6), showing that models tend to predict simpler cascades with fewer relations and programs, often succeeding only when a functionally equivalent solution exists.

### v4

@@ 3.1.3 Enforcing Complexity Constraints
We control the complexity of cascades of programs $\vec{p}$ using rejection sampling guided by the classifiers developed in Section B. Complexity is balanced along two dimensions: (1) the cascade length $L=|\vec{p}|$, and (2) the relation types present between programs in $\vec{p}$. The second dimension is encoded as a binary category string per cascade $\vec{p}$, $c_{\vec{p}}\in\{0,1\}^{4}$. Each digit is a binary indicator denoting the presence or absence of feeding, bleeding, counter-feeding, and counter-bleeding relations. Instances are balanced across all 16 possible category strings, from $c_{\vec{p}}=0000$ (no relations present, arbitrary ordering possible) to $c_{\vec{p}}=1111$ (all relations present, ordering highly constrained).
Computing Instance Complexity. We hypothesize that the difficulty of a PBE problem $\langle\vec{\imath},\vec{p},\vec{o}\rangle$ is governed by (1) cascade length and (2) the types of relations between program pairs in $\vec{p}$ (relations formulated by phonologists and historical linguists; see Kiparsky (1968, 1971)). One of our key contributions is provably correct, automatic, symbolic classification of these relations for a given pair of programs (Appendix B). For a pair of programs ($p_{i}$$p_{j}$) in $\vec{p}$, where $p_{i}$ is applied before $p_{j}$, the possible relations are:
Feeding (F($p_{i}$, $p_{j}$)): $p_{i}$ creates substrings that enable $p_{j}$ to apply.
Bleeding (B($p_{i}$, $p_{j}$)): $p_{i}$ removes substrings required by $p_{j}$.
@@
Counter-relations need not be stored explicitly, since $CF(p_{i},p_{j})$ and $CB(p_{i},p_{j})$ are implied if $p_{j}$ precedes $p_{i}$ and $F(p_{j},p_{i})$ and $B(p_{j},p_{i})$, respectively. These relations can be visualized using a Directed Acyclic Graph (DAG), as illustrated in Figure 1(b). We show both the full DAG with all relations and a simplified DAG that indirectly encodes counter-relations via this symmetry.
Rejection Sampling to Control Complexity. To enforce balanced complexity, we use rejection sampling. Each PBE instance is first assigned a relation-type category string $c_{\vec{p}}$, and data is generated to approximate balance across the $2^{4}=16$ categories, as well as across cascade lengths between $L_{min}$ and $L_{max}$. These constraints can conflict, since short cascades (e.g., $L=2$) rarely realize higher-order relation types, making some categories unattainable. Following prior work on sound law induction (Naik et al., 2025a), we retain random sampling rather than deterministic generation or retrieval to promote diversity and avoid contamination, at the cost of high rejection rates. We introduce a patience parameter $\tau$, denoting the number of sampling steps under both constraints; once $\tau$ (=100,000=100{,}000) is exceeded, we also accept instances satisfying just one constraint. Early experiments prioritized relation-type balance, but later analysis (Appendix F.7, F.8) identified cascade length as a stronger predictor of difficulty, leading us to relax relation-type constraints after excceeding $\tau$ steps. We show the effect of varying $\tau$ on relation-type balance and generation efficiency (Appendix F.6).
@@ 3.2 Problem Solver
Prompting Strategy: Once a benchmark snapshot $\mathcal{D}$ is generated, we evaluate each LLM $M$ by prompting it for multi-step PBE (second from bottom in Fig 1(a); prompt in Appendix D.3.1) and program reordering (bottom in Fig 1(a); prompt in Appendix D.3.2).
For multi-step PBE, the prompt enforces the following constraints: each program must be a Python replace function; both arguments $\alpha_{k}$ and $\beta_{k}$ must be strings with $|\alpha_{k}|,|\beta_{k}|\leq s_{max}$; $\alpha_{k}$ must be non-empty; the cascade may contain at most $L_{max}$ programs (matching the longest ground-truth cascade); and outputs must follow a strict Markdown format for reliable extraction. If a predicted cascade $\hat{\vec{p}}$ exceeds $L_{max}$ programs, only the first $L_{max}$ are evaluated; violations of any other constraint result in replacement with an identity program $p^{I}$, which leaves inputs unchanged. Given $\hat{\vec{p}}$, predicted outputs are computed as $\hat{\vec{o}}=\hat{\vec{p}}(\vec{\imath})$. Details of program extraction from LLM responses are provided in Appendix D.2.
For the program reordering task, the prompt defines bleeding and feeding relations, provides the ground-truth program cascade along with the inputs and outputs, and instructs the LLM to output only a fenced JSON block containing a permutation of indices (e.g., $[2,1,3]$) that recovers the original program order. This task evaluates the model’s ability to reason about ordering constraints induced by BFCC relations between programs.
Scaling Solution Search. We investigate two solution-search scaling strategies, motivated by evidence of reasoning collapse and out-of-token failures in (Shojaee et al., 2025) and subsequent discussion (Lawsen, 2025): (1) sampling-budget (Li and Ellis, 2024) and (2) test-time thinking-budget scaling (Muennighoff et al., 2025). In the first, the LLM produces $K$ candidate solutions, and any candidate consistent with the input–output examples is accepted; if none succeed, we select the candidate with the highest edit-similarity reward (Section 4.3). In the second, inspired by Muennighoff et al. (2025), we vary a thinking budget of $N$ tokens. However, because gpt-oss-120b does not reliably respect this constraint, even when forced to emit chain-of-thought termination tokens (Appendix E.3), we instead control the overall maximum sequence length.
@@ 4 Experiments
4.1 Dataset Creation
Using the problem proposer described in Section 3.1, we construct three synthetic benchmarks.
PBEBench-Lite: A simplified dataset with cascades of length 2 to 5 and 5 input–output pairs per instance, synthesized as $\mathcal{D}(n=5,\Sigma=\Sigma_{\text{lite}},L_{min}=2,L_{max}=5,l_{min}=2,l_{max}=6,s_{min}=1,s_{max}=3,D=1008)$, with exactly 63 examples per relation category. The alphabet is restricted $\Sigma_{\text{lite}}=\{a,\dots,k,u,v,w,x,y,z\}$ (lowercase letters excluding lltt).
PBEBench-Lite-Perm: A program reordering dataset with 919 instances created from PBEBench-Lite using the swapping strategy (Section 3.1).
PBEBench: a larger dataset with 50 input-output pairs per instance, synthesized as $\mathcal{D}(n=50,\Sigma,L_{min}=2,L_{max}=20,l_{min}=2,l_{max}=6,s_{min}=1,s_{max}=3,D=1216)$, with 64 instances for each cascade length. Here $\Sigma=\{a,\dots,z,A,\dots,Z\}$ (full alphabet in both cases).
Real-SLI: We also evaluate all open-source LLMs on real SLI data drawn from 6 proto–attested language pairs, described in more detail in Appendix C.3. To assess the predictive power of PBEBench for SLI performance, we additionally compare it against other inductive reasoning benchmarks, including CLUTRR Sinha et al. (2019) and SLR-Bench Helff et al. (2025). Some other datasets generated for some ablations are described in Appendix C.2.
@@ 4.2 Models Evaluated
We evaluate a broad range of state-of-the-art LLMs spanning multiple model families and architectures. We cover models from leading developers including OpenAI (GPT, o-series, gpt-oss), Anthropic (Claude series), Google (Gemini), Qwen (Qwen2.5, Qwen3, QwQ), DeepSeek (R1-Distill-Qwen), Mistral (Codestral). The models differ in scale, reasoning specialization (thinking vs. non-thinking), architectural choices (dense vs. MoE), and source availability (open vs. closed). A full list of models with their attributes is provided in Table 4, covering the breadth of model families and capabilities.
@@ 4.3 Evaluation Metrics
For the multi-step PBE task, since a given input vector $\vec{\imath}$ may be mapped to the output vector $\vec{o}$ by multiple program cascades $\vec{p}$, we evaluate model predictions using metrics based on functional equivalence, following prior work in programming by example Li and Ellis (2024) and historical linguistics Hoenigswald (1960). We execute the model-generated cascade $\hat{\vec{p}}$ on the inputs and compare the predicted outputs $\hat{\vec{o}}=\hat{\vec{p}}(\vec{\imath})$ with $\vec{o}$.
We use two main metrics: a coarse-grained solve rate (Pass@1) and a fine-grained normalized edit similarity (Edit_Sim):
$\mathrm{pass@1}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{\imath}\in\mathcal{D}}\mathrm{1}_{\hat{\vec{p}}(\vec{\imath})=\vec{o}}$
$\mathrm{Edit\_Sim}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{\imath}\in\mathcal{D}}1-\frac{\mathrm{dist}(\hat{\vec{p}}(\vec{\imath}),\vec{o})}{\mathrm{dist}(\vec{\imath},\vec{o})},$
where $\mathrm{dist}$ denotes the summed Levenshtein edit distance over corresponding strings. We additionally report the Valid_Rate, the proportion of generated programs that satisfy all prompt constraints and Complexity that sums up the length of the substrings $\alpha_{k}$ and $\beta_{k}$ in the model predicted program cascade and averages it across PBE instances. For the program reordering task, we evaluate accuracy (Acc) by applying the predicted permutation $\hat{\sigma}$ to the permuted program sequence and checking whether $\hat{\sigma}(\sigma(\vec{p}))(\vec{\imath})=\vec{o}$ (Section C.4). We also report unique accuracy on instances with a single valid solution.
Model Acc UAcc
Codestral-22B 34.3 51.2
Qwen2.5-32B-Instruct 31 39.3
Qwen2.5-Coder-32B-Instruct 34.7 49.6
Qwen3-32B 36 43
Qwen3-Coder-30B-A3B-Instruct ◼ 44.1 63.2
DeepSeek-R1-Distill-Qwen-32B ★ 75.5 92.1
Qwen3-30B-A3B ★◼ 68 91.3
QwQ-32B ★ 74.9 94.6
Qwen3-32B (with CoT) ★ 77.9 93.8
gpt-oss-20b ★◼ 86.3 92.6
gpt-oss-120b ★◼ 97.5 98.8
Claude-4-Sonnet ✸ 80.2 91.3
Claude-4.5-Sonnet ✸ 85.1 92.1
Claude-4-Sonnet (Thinking) ★✸ 91.6 97.5
Claude-4.5-Sonnet (Thinking) ★✸ 97.5 99.2
o3-mini ★✸ 63.4 83.7
o4-mini ★✸ 80.2 97.5
Gemini 2.5 Flash ★✸ 92.8 89.7
GPT-5 ★✸ 99.7 99.6
Table 1: PBEBench-Lite Program Reordering: We compute accuracy (Acc) based on functional correctness of the unscrambled program sequence on the full data and the unique solution subset (UAcc). ◼ indicates a mixture-of-experts (or MoE) model. ★ indicates a reasoning model. ✸ indicates unknown architecture.
Model Pass@1 Edit Sim Complexity Valid Rate
Codestral-22B 1.1 -1.1 15.4 82.5
Qwen2.5-32B-Instruct 3 12.5 12.5 82.9
Qwen2.5-Coder-32B-Instruct 4.1 18.8 12.67 68.9
Qwen3-32B 1.8 9.6 14.89 76.5
Qwen3-Coder-30B-A3B-Instruct ◼ 3.8 8.6 3.25 81.3
DeepSeek-R1-Distill-Qwen-32B ★ 22.4 34.9 7.43 87.1
Qwen3-30B-A3B ★◼ 28.9 33.6 4 99.1
QwQ-32B ★ 36 40.9 4.72 94.9
Qwen3-32B (with CoT) ★ 41.9 50.3 6.57 96.8
gpt-oss-20b ★◼ 40.6 46.2 6.96 99
gpt-oss-120b ★◼ 62.5 69.9 10.93 92.5
Claude-3.5-Sonnet ✸ 18.5 44.3 11.87 82.1
Claude-3.7-Sonnet ✸ 23.2 50 12.05 84.1
Claude-4-Sonnet ✸ 29.7 58.7 13.35 77.2
Claude-3.7-Sonnet (Thinking) ★✸ 36.6 61.5 12.94 0.819
Claude-4-Sonnet (Thinking) ★✸ 35.8 60.8 13.71 78.2
Claude-4-Opus (Thinking) ★✸ 53.9 75.2 13.44 85.6
o3-mini ★✸ 19.6 19.8 1.65 99.5
o4-mini ★✸ 36.3 37.4 4.08 97.3
Gemini 2.5 Flash ★✸ 58.6 65.6 8.36 79
GPT-5 ★✸ 72.4 76.5 10.58 92.9
Table 2: PBEBench-Lite Performance: We compute the Pass@1 and Edit_Sim as the coarse and fine-grained evaluation, respectively, for each model. ◼ indicates a mixture-of-experts (or MoE) model. ★ indicates a reasoning model. ✸ indicates unknown architecture. * indicates evaluated on 20% of the dataset due to cost.
@@ 5.1 PBEBench-Lite Performance
Table 2 reports Pass@1, Edit_Sim, Valid_Rate (expressed as percentages), and Complexity for all models (Section 4.2). Hyperparameters used in model evaluations are listed in Table 8. Claude-4 Opus (Thinking) was evaluated on only 20% of the data due to high API costs (run costs are reported in Appendix D.6). Reasoning models, both open and closed-source, consistently outperform non-reasoning models, with a larger gap among open-source models. The top-performing models are gpt-oss-120b (open) and GPT-5 (closed). Surprisingly, Qwen3-30B-A3B and o3-mini produce the simplest programs yet perform poorly, suggesting underthinking, while Claude-4-Sonnet and Qwen3-32B without CoT produce the most complex programs, indicative of overthinking. The ground-truth programs have a mean complexity of 11.63, indicating that even the best-performing models, GPT-5 and gpt-oss-120b, generate simpler programs than the ground truth. Factorial analysis of QwQ-32B and GPT-5 identifies cascade length as the strongest negative predictor of Pass@1. We further find that feeding and bleeding decrease Pass@1, whereas counter-feeding and counter-bleeding have no significant effect (Appendix F.7). Finally, confusion-matrix analyses of predicted versus ground-truth cascade lengths and relation types (Appendix F.10, F.11) show that models favor simpler cascades with fewer relations and programs, typically succeeding only when a functionally equivalent solution exists.
5.2 PBEBench-Lite-Perm Performance
Table 1 reports Acc and UAcc for the PBEBench-Lite-Perm program reordering dataset. We could not run Claude-3.5-Sonnet and Claude-3.7-Sonnet due to deprecation, and Claude-4-Opus due to budget constraints, and therefore evaluate Claude-4.5-Sonnet with and without reasoning. While LLMs may sometimes bypass BFCC relations, this task requires explicit reasoning to recover correct ordering, with the unique-solution subset admitting only one valid order. Using PBEBench-Lite with short cascades (2–5) allows some models to enumerate all orderings, simplifying the task relative to full multi-step PBE. GPT-5 nearly solves the task with 99.7% accuracy, whereas Qwen2.5-32B-Instruct performs worst at 31%. Logistic regression analysis (Table 17) reveals that LLMs predictably struggle more with longer cascades and counter relations, like counter-feeding. Finally, UAcc is generally higher than Acc, explained by the fact that 182 of 242 unique-solution instances have cascade length 2, where random guessing already yields 50% accuracy. We also present a qualitative analysis of the various solution strategies adopted by the evaluated models in Appendix F.9.